A) Explain why the higher the MPC, the higher the RDGP: Answer:

B)With a MPC of 68%, what is the Multiplier (M)? Answer:

C)What does Keynes think the answer to recessionary gap is based on this chapter? Answer:

D) There are income taxes and autonomous taxes. Which one would classical economists believe should be primary and which one would Keynesians choose? Explain. Answer:
Answer #1

A.

A higher MPC means, more of the income earned will be consumed and it will lead to increase in aggregate demand (AD).  Further, a higher MPC, will have higher multiplier in the economy, making degree of increase in AD to be bigger. As a result, a bigger right shift in AD takes place and real GDP increase with a bigger margin.

B.

Multiplier = 1/(1-MPC) = 1/(1-.68)  

Multiplier = 3.125

C.

The answer to recessionary gap is to increase the government spending and or decrease the tax rate. These two initiatives, will increase the AD and firms will cater the Ad by increasing their SRAS. It will create jobs and more consumption and rise in AD takes place. It will make a chain of positive economic activities and economy will recover form the recessionary gap.

D.

Classical economist will make autonomous taxes should be primary, and Keynes will give focus upon income taxes to help economy achieve the full employment level.